What is thedimensional formula of work . Please derive it 

Grade:12

2 Answers

Vikas TU
4 years ago
Dear student 
Work is defined is the force displaced through a distance.
i .e., Work = Force × distance
= (mass× acceleration) × distance
= (M× L× T^-2) ×L
=M×L^2×T^-2.
Hope this helps 
Good luck 
Arun
4 years ago
The formula for work is 
Force* displacement 
 W=Force * displacement 
 = Mass* acceleration* displacement 
  =M*(velocity/time)*displacement 
  = M1L2T-2 
Here force is regarded as the product of mass and acceleration. 
 M represents mass; L is regarded as length while T is considered as time. 
The dimensional formula is derived from the conventional formul